A certain amount of money doubles after 9 years. Determine its effective interest.
1
Expert's answer
2020-09-21T09:06:35-0400

The precise formula for calculating the exact doubling time for an investment earning a compounded interest rate of r% per period is as follows:

"T=\frac{ln2}{ln(1+\frac{r}{100})}≈\frac{72}{r}"

where: T=Time to double;

ln⁡=Natural log function;

r=interest rate per period

≈Approximately equal to

"9=\frac{72}{r}"

"r=\frac{72}{9}=8%"